OpenStudy (anonymous):

Solve x2 + 10x + 24 = 0 by completing the square. Show your work for full credit.

OpenStudy (anonymous):

\[(x^2+10x+24=0\] \[x^2+10x=-24\] \[(x+5)^2=-24+25=1\] \[x+5=\pm1\] \[x=1-5, x=-1-5\] \[x=-4,x=-6\]

OpenStudy (amistre64):

x2 + 10x + 24 = 0 (x2 + 10x) + 24 = 0 (x) (x + 10) + 24 = 0 (x+5) (x+5) + 24-5^2 = 0 etc